Kalin walks at a steady rate of 3 2/3 miles per hour. The beach is 4 1/4 miles from home. How long will it take him to walk from his home to the beach and back to his home? We know he has to walk a total of 8 1/2 miles but cant figure out the rest.
Please Help

Answer by rfer(16322)

(17/2)/(11/3)=
17/2*3/11=51/22=2.3 hrs
-----------------------
8.5/3.667=2.3 hrs or 2 hrs 18 min